Circular No. 9015

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 WZ_{96} (LINEAR) An apparently asteroidal object that was discovered by the LINEAR

survey in late November and again in late December (when the Minor Planet Center linked the two sets of observations), and designated 2008 WZ_{96} (discovery observations tabulated below; cf. MPS 267760, 270506), has been found to show cometary appearance by CCD astrometrists elsewhere. P. C. Sherrod (Conway, AR, U.S.A., 0.51-m f/4.3 astrograph) reports that images taken on Jan. 20.05–20.07 UT show a round, small, uniform nebulosity of total mag ∼ 17.7 surrounding a slight central condensation of mag ∼ 18.5; he adds that images taken over the previous two weeks had also shown suggestions of a diffuse nature. J. Y. Young (Table Mountain 0.61-m f/16 Cassegrain reflector) writes that his images from Jan. 29.1 show a round coma of diameter 6′′ with only a moderate condensation and no tail visible even in longer exposures. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m f/10 Schmidt-Cassegrain reflector) notes that twelve stacked 60-s exposures from Jan. 30.1 show a round 12′′ coma with a stellar nucleus condensation and no tail.

2008 UT α_{2000} δ_{2000} Mag. Observer
Nov. 30.06444 22^{h}24^{m}11.^{s}08 −16^{o}41′49′′.6 18.3 LINEAR
Dec. 28.06701 23 21 01.03 − 8 43 02.4 18.1  _{”}

Additional astrometry, the following orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-B73. Marsden notes that the comet passed 0.36 AU from Jupiter in Apr. 2006.

Epoch = 2009 Jan. 9.0 TT

T = 2009 Jan. 23.9424 TT  ω = 337.^{o}6385
e = 0.509060  Ω = 59.1168 }2000.0
q = 1.646014 AU  i = 6.9574

a = 3.352781 AU n^{o} = 0.1605448 P = 6.14 years

COMET P/2009 B1 (BOATTINI) Additional astrometry (including pre-discovery Catalina observations

from 2008 Nov. 18) of this comet (cf. IAUC 9013) and the following revised orbital elements by B. G. Marsden were published on MPEC 2009-B50: Epoch = 2009 Feb. 18.0 TT, T = 2009 Feb. 6.2408 TT, e = 0.637146, q = 2.426329 AU, ω = 128^{o}.6057, Ω = 297^{o}.4434, i = 22^{o}.2273 (equinox 2000.0), P = 17.29 years.

2009 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green